# Understanding Common Ground Misalignment in Goal-Oriented Dialog: A Case-Study with Ubuntu Chat Logs

### v1

@@ Abstract
While it is commonly accepted that maintaining common ground plays a role in conversational success, little prior research exists connecting conversational grounding to success in task-oriented conversations. We study failures of grounding in the Ubuntu IRC dataset, where participants use text-only communication to resolve technical issues. We find that disruptions in conversational flow often stem from a misalignment in common ground, driven by a divergence in beliefs and assumptions held by participants. These disruptions, which we call conversational friction, significantly correlate with task success. We find that although LLMs can identify overt cases of conversational friction, they struggle with subtler and more context-dependent instances requiring pragmatic or domain-specific reasoning.
@@ 2 Background
wThe conversational common ground (CG) is a body of statements treated as mutual knowledge among participants Stalnaker (1978). It guides both how speakers choose their utterances and how they want them to be interpreted Stalnaker (2002) [^6] Even before Stalnaker, Paul Grice mentioned propositions having common ground status in his William James lectures Stalnaker (2002). For a thorough discussion of common ground in linguistics, see Geurts (2024).. Subsequently, Clark and Brennan (1991) define common ground as a collection of mutual knowledge, beliefs, and assumptions that humans maintain collaboratively through the process of grounding. [^7] We focus only on discourse-theoretic grounding and do not delve into symbol grounding Harnad (1990), as exemplified in mapping a linguistic concept to a visual scene (see Cohen et al. (2024) for a survey of methodologies for robotic language grounding); however, we embrace the conceptual relationship between both types of grounding, as described in Chandu et al. (2021).
@@
Typical conversations in our dataset (e.g. in Table 3) involve two participants (an asker and a helper) collaboratively attempting to solve a Linux bug over a text channel. This consists of several steps—the asker must describe their issue (often with insufficient knowledge of Linux), and the helper must understand their goal to propose a solution. This makes the setting well-positioned for studying friction and grounding.
@@
The Ubuntu Dialog Corpus Lowe et al. (2015) contains conversations scraped from the #Ubuntu IRC channel, where users primarily discuss features, issues and bugs related to the Ubuntu operating system. This requires disentangling conversations from a single message stream. Kummerfeld et al. (2019) found that the disentanglement strategy originally used had a high error rate, and release a cleaner version. We use a sample of 200 two-person conversations from this cleaner corpus, upsampling longer conversations to study diverse behavior (Table 1). We refer to this subset as Ubuntu-CG (Common Ground).
@@
1 (No Progress) 33.05 (25.84) 64.40 (38/59) 2.39
2 (Some Progress) 44.22 (25.49) 64.19 (52/81) 2.08
3 (Success) 40.3 (28.56) 53.33 (32/60) 2.09
@@ 3.1 Annotating Conversational Friction
Three computer science undergraduates familiar with Linux annotate conversations by (1) identifying turn intervals which exhibit friction along with explanations and (2) judging the success of overall conversation on a three-point scale with respect to the conversational goal. Annotators were paid \$18/hr to annotate 200 conversations totaling 7950 turns, taking over 80 hours to complete. Since conversations date back to over a decade ago, they often contain antiquated terms or references that annotators were unfamiliar with. To mitigate this, we provide explanations generated by gpt-4o OpenAI (2024) of technical terms in dialog turns. For example, the model-generated elaboration in Table 9 (Row 1) explains that “dapper” and “feisty” refer to Ubuntu versions 6.06 and 7.04. We make these elaborations available to various models in our computational experiments as well.
@@
In addition to friction, annotators assess how successful participants were in solving the issue by scoring the conversation on a three-point scale. A score of 1 denotes that the conversation was not helpful to the asker at all, and no progress was made; a score of 2 denotes some progress towards solving or diagnosing the issue, and a score of 3 indicates that the issue was solved. In cases where experienced helpers propose alternate solutions, success is measured by progress towards this new goal. Table 2 shows overall statistics, and the instructions for friction and success annotation can be found in the Appendix A.6. We obtain an agreement of $\alpha=0.58$ on success annotation as measured by Krippendorff’s Alpha Castro (2017).
@@ 3.2 Measuring Friction Agreement
We consider a second setting that rewards the degree of overlap with the gold interval. We first match each instance of friction with the predicted instance with the highest overlap, ensuring a one-to-one mapping between a predicted and gold friction interval. For each matched interval, we compute the Jaccard similarity between the two intervals, resulting in higher scores for predictions that better align with human-annotated instances of friction and penalizing predicting multiple short or overtly long instances. [^11] This is similar in spirit to methods discussed in Ortmann (2022), adapted for our task. A perfect score indicates an exact overlap between predicted and gold intervals. We use these same two settings to compute model performance (Table 6). Table 4 shows agreement between pairs of annotators.
@@ 3.3 Annotating for Grounding Acts
We sample 70 conversations containing 152 instances of friction to study the effects of grounding on task success. 21 conversations received a success score of 1 (No Progress), 26 received a score of 2 (Some Progress), and 23 received a score of 3 (Success). Since conversations with friction tend to be longer, this sample has a higher average length than our dataset overall. Two authors annotated each friction instance in this subset for the presence or absence of RequestRepair and Repairacts, obtaining inter-rater scores of $0.69$ on RequestRepair , and $0.63$ on Repair , measured using Cohen’s Kappa Cohen (1960).
@@ 4 Analysis of Grounding in Ubuntu-CG
In Ubuntu-CG, 61% percent of conversations contained friction. In contrast, of conversations where the helper succeeded in solving the asker’s issue (receiving a score of 3), only 53.33% contained friction (Table 2). Conversations where participants make some progress or succeed contain less friction on average as compared to conversations where they did not make any progress, as the former exhibits some amount of grounding effort by the participants (Table 2, Column 4). This is further supported by the proportion of unaddressed repair efforts (Table 5, Column 4).
@@
While no clear pattern emerges in conversation length varies with task success (Table 2), the mean length of a conversation containing friction is 49 (median 50), as compared to an average of 20.08 (median 12) of those without friction. Compared to the overall mean length of the dataset 40.56 (median 33), it is plausible that conversational friction and repair through the process of grounding contributes to the increased number of turns it takes to complete the conversation.
@@ 5 Can LLMs Identify Conversational Friction?
gpt-4o 40.10 71.43 51.36 17.00 30.28 21.77 495
gpt-4o w/ Elab. 41.83 65.18 50.96 18.03 28.09 21.96 435
gpt-4o-mini 38.37 44.20 41.08 16.36 18.84 17.51 316
gpt-4o-mini w/ Elab. 32.72 47.77 38.84 16.65 24.30 19.76 392
Llama-3.1-8b-Instruct 17.94 81.70 29.42 7.08 32.25 11.62 1282
Llama-3.1-8b-Instruct w/ Elab. 19.08 87.05 31.30 7.58 34.60 12.44 1253
Llama-3.1-70b-Instruct 26.55 82.14 40.13 10.38 32.12 15.69 857
Llama-3.1-70b-Instruct w/ Elab. 21.35 75.00 33.23 8.81 30.94 13.71 959
distilroberta-base ($k=3$, finetuned) 23.30 49.54 31.53 8.86 18.77 11.98 -
distilroberta-base ($k=5$, finetuned) 20.86 47.04 28.81 9.50 21.15 13.06 -
Table 6: Precision, Recall, and F1 scores of different models on detecting friction. #Predictions refer to the total number of instances of conversational friction found by each model. For reference, annotators identified 266 instances in total. gpt-4o with Elaboration of technical terms (Sec 3.1) performed best across all models.
@@ 5.1 Experimental Setup
We explore a baseline setting in which we finetune a small encoder-based model distilroberta-base Sanh et al. (2019) on five randomly-split folds of Ubuntu-CG (Appendix A.2). Given an excerpt of a conversation consisting of a target turn $t$ and a context window of $k$ turns before and after $t$, the model is trained to predict whether $t$ is part of an annotated instance of friction. Finetuning a lighter-weight model allows us to understand the extent to which friction is identifiable from surface features.
@@
Given a full conversation, we prompt several larger decoder-only LLMs to output a list of turn intervals exhibiting friction (Prompt A.4 in Appendix). [^12] We experimented with several prompting strategies such as adding random exemplars, self-consistency, and chain-of-thought reasoning, but found that they did not beat the F1 scores obtained simply by asking the model to detect friction windows along with brief explanations of why a dialog window represents friction. For each predicted turn interval, the model must provide a brief explanation for the cause of friction. We also include a setting where we provide models with the elaborations of technical terminology as outlined in §3.1 (“w Elab”).
@@
We evaluate models in the Friction Found and Friction Overlap settings 3.2. While Friction Found allows models like Llama-3.1-8b-Instruct Touvron et al. (2023) to obtain high recall scores by over-predicting friction intervals, Friction Overlap penalizes this behavior. For all experimental settings, we set temperature to $0.01$.
@@ 5.2 Results
The F1 scores of our baseline encoder-only models (Table 6) in two settings (context window $k=3$ and $k=5$) give us an estimate of the degree to which friction is identifiable from shallower, local features as opposed to more complex, contextual, and implicit cases of friction like subtle clarification questions or other pragmatic phenomena. Error analysis reveals that instances predicted correctly by our baselines often contained explicit markers of friction such as a user expressing frustration or dissatisfaction. While these baselines were competitive with zero-shot Llama-3.1-8b, it is difficult to interpret the results beyond error analysis since we do not have access to explanations. However, it does indicate that our dataset contains learnable signals.
Under both evaluation settings, gpt-4o or gpt-4o with elaborations obtained the highest F1 score. llama models have the highest recall, which is balanced by their low precision. Under the stricter “span” setting, gpt-4o with elaboration had the highest F1 score. We use this setting for all further error analysis and ablations. All models over-predict friction intervals (see column #Predictions in Table 6).
@@
Explaining technical terms with gpt-4o helped our human annotators better understand the flow of information in a conversation. However, in the relaxed evaluation setting (Friction Found), adding elaborations does not improve prediction scores of models, though Friction Overlap yields stronger performance across all models except Llama-3.1-70b-Instruct. For Llama-3.1-8b-Instruct and gpt-4o-mini, adding elaborations improves both precision and recall. This may be due to elaborations “sharpening” the predicted intervals.
@@
We evaluate the binary prediction task with Cohen’s $\kappa$, framing it as inter-rater agreement between models and humans. Models’ over-prediction of friction intervals persist in the conversation level as well (Table 7). Predictions on task success, on the other hand, is highly correlated with annotator ratings of success.
@@ 6 Error Analysis
We now investigate the successes and failures of our strongest performing setting, gpt-4o with elaborations.
@@
As a conversation proceeds, detecting friction requires a deeper understanding of preceding turns. To explore whether the position of friction impacts model accuracy, we stratify our results by conversational depth and calculate the relative depth of each instance of friction as the ratio of the first turn of the friction interval to the conversation length multiplied by 100. The mean relative depth of a detected instance of friction ($35.19$) is significantly smaller than the mean relative depth of a detected instance ($49.62$), according to an independent t-test ($p<0.01$). This indicates that models struggle with taking longer context into account while determining whether participants’ versions of common ground are misaligned.
@@
Models, particularly gpt-4o with elaboration, are more likely to correctly identify friction when an explicit request for conversational repair is present. Specifically, 77.22% of detected frictions involved an explicit RequestRepair, compared to $64.81$% of frictions that went undetected ($p<0.05$). This highlights the tendency of models to rely on overt cues that signal a common ground misalignment.
Consider the conversation in Table 8. A’s utterance “how about nmap” (Turn 22) is not introducing nmap as an option; but following up on B’s earlier suggestion in Turn 21 by asking how nmap can be used to solve the issue. B reveals that they did not understand this interpretation through their response in Turn 23 (“yeah, i said nmap,”), prompting A to issue a Repair act. We hypothesize that this unconventional way of issuing a Repair (through a question) without an explicit RequestRepair results in an undetected conversational friction.
@@ 24 A (asker) I mean how do I use nmap to find that out?
Most model explanations accurately captured the cause of friction, with $57.81$% instances annotated as equivalent and $34.37$% as somewhat similar. Only $7.8$% explanations did not point out the cause of friction at all. Echoing our earlier findings, models struggle to pinpoint the cause of friction even when they identified the window correctly. For example, in Figure 2, Row 1, A mistakenly assumes that B’s suggestion for a command (Turn 25) includes the keyword “try”, leading to the error “try command not found” (Turn 32) after which B repeats the dmesg command removing “try” at the beginning. LLMs must be able to pinpoint causes of friction to issue repairs that address the friction directly, a crucial ability in settings where LLMs are used for dispute resolution Tan et al. (2024).

### v2

@@ Abstract
While it is commonly accepted that maintaining common ground plays a role in conversational success, little prior research exists connecting conversational grounding to success in task-oriented conversations. We study failures of grounding in the Ubuntu IRC dataset, where participants use text-only communication to resolve technical issues. We find that disruptions in conversational flow often stem from a misalignment in common ground, driven by a divergence in beliefs and assumptions held by participants. These disruptions, which we call conversational friction, significantly correlate with task success. While LLMs can identify overt cases of conversational friction, they struggle with subtler and more context-dependent instances that require pragmatic or domain-specific reasoning.
@@ 2 Background
The conversational common ground (CG) is a body of statements treated as mutual knowledge among participants Stalnaker (1978). It guides both how speakers choose their utterances and how they want them to be interpreted Stalnaker (2002) [^6] Even before Stalnaker, Paul Grice mentioned propositions having common ground status in his William James lectures Stalnaker (2002). For a thorough discussion of common ground in linguistics, see Geurts (2024).. Subsequently, Clark and Brennan (1991) define common ground as a collection of mutual knowledge, beliefs, and assumptions that humans maintain collaboratively through the process of grounding. [^7] We focus only on discourse-theoretic grounding and do not delve into symbol grounding Harnad (1990), as exemplified in mapping a linguistic concept to a visual scene (see Cohen et al. (2024) for a survey of methodologies for robotic language grounding); however, we embrace the conceptual relationship between both types of grounding, as described in Chandu et al. (2021).
@@
Typical conversations in our dataset (e.g., in Table 3) involve two participants (an asker and a helper) collaboratively attempting to solve a Linux bug over a text channel. This consists of several steps—the asker must describe their issue (often with insufficient knowledge of Linux), and the helper must understand their goal to propose a solution. This makes the setting well-positioned for studying friction and grounding.
@@
The Ubuntu Dialog Corpus Lowe et al. (2015) contains conversations scraped from the #Ubuntu IRC channel, where users primarily discuss features, issues, and bugs related to the Ubuntu operating system. This requires disentangling conversations from a single message stream. Kummerfeld et al. (2019) found that the disentanglement strategy originally used had a high error rate, and released a cleaner version. We use a sample of 200 two-person conversations from this cleaner corpus, upsampling longer conversations to study diverse behavior (Table 1). We refer to this subset as Ubuntu-CG (Common Ground).
@@
1 (No Progress) 31.90 (24.99) 57.60 (30/52) 2.43
2 (Some Progress) 43.86 (25.73) 55.05 (49/89) 2.06
3 (Success) 40.45 (28.84) 50.84 (30/59) 2.13
@@ 3.1 Annotating Conversational Friction
Three computer science undergraduates familiar with Linux annotate conversations by (1) identifying turn intervals which exhibit friction along with explanations and (2) judging the success of overall conversation on a three-point scale with respect to the conversational goal. Annotators were paid \$18/hr to annotate 200 conversations totaling 7950 turns, taking over 80 hours to complete. Since conversations date back to over a decade ago, they often contain antiquated terms or references that annotators were unfamiliar with. To mitigate this, we provide explanations generated by gpt-4o OpenAI (2024) of technical terms in dialog turns. For example, the model-generated elaboration in Table 9 (Row 1) in Appendix explains that “dapper” and “feisty” refer to Ubuntu versions 6.06 and 7.04. We make these elaborations available to various models in our computational experiments as well.
@@
In addition to friction, annotators assess how successful participants were in solving the issue by scoring the conversation on a three-point scale. A score of 1 denotes that the conversation was not helpful to the asker at all, and no progress was made; a score of 2 denotes some progress towards solving or diagnosing the issue, and a score of 3 indicates that the issue was solved. In cases where experienced helpers propose alternate solutions, success is measured by progress towards this new goal. Table 2 shows the overall statistics, and instructions for friction and success annotation can be found in the Appendix A.6. We obtain an agreement of $\alpha=0.58$ on success annotation as measured by Krippendorff’s Alpha Castro (2017).
@@ 3.2 Measuring Friction Agreement
We consider a second setting that rewards the degree of overlap with the gold interval. We first match each instance of friction with the predicted instance with the highest overlap, ensuring a one-to-one mapping between a predicted and gold friction interval. For each matched interval, we compute the Jaccard similarity between the two intervals, resulting in higher scores for predictions that better align with human-annotated instances of friction and penalizing predicting multiple short or overtly long instances. [^11] This is similar in spirit to methods discussed in Ortmann (2022), adapted for our task. A perfect score indicates an exact overlap between predicted and gold intervals. We use these same two settings to compute model performance (Table 6). Table 4 shows agreement between pairs of annotators. Given their high agreement, A1 and A2 annotated 80% of the final dataset, while A3 annotated 20%.
@@ 3.3 Annotating for Grounding Acts
We sample 70 conversations containing 152 instances of friction to study the effects of grounding on task success. 21 conversations received a success score of 1 (No Progress), 26 received a score of 2 (Some Progress), and 23 received a score of 3 (Success). Since conversations with friction tend to be longer, this sample has a higher average length than our dataset overall. Two authors annotated each friction instance in this subset for the presence or absence of RequestRepair and Repairacts, obtaining inter-rater scores of 0.690.690.69 on RequestRepair , and 0.630.630.63 on Repair , measured using Cohen’s Kappa Cohen (1960).
@@ 4 Analysis of Grounding in Ubuntu-CG
In Ubuntu-CG, 61% percent of conversations contained friction. In contrast, of conversations where the helper succeeded in solving the asker’s issue (receiving a score of 3), only 54.5% contained friction (Table 2). Conversations where participants make some progress or succeed contain less friction on average as compared to conversations where they did not make any progress, as the former exhibits some amount of grounding effort by the participants (Table 2, Column 4). This is further supported by the proportion of unaddressed repair efforts (Table 5, Column 4).
@@
While conversation length shows no clear pattern with task success (Table 2), the mean length of a conversation containing friction is 49 (median 55), as compared to an average of 29 (median 22) of those without friction. Compared to the overall mean length of the dataset 40.56 (median 33), it is plausible that conversational friction and repair through the process of grounding contribute to the increased number of turns it takes to complete the conversation.
@@ 5 Can LLMs Identify Conversational Friction?
gpt-4o 31.50 43.69 34.01 13.50 18.74 14.61 495
gpt-4o w/ Elab. 31.63 37.46 32.22 13.54 16.59 14.00 435
gpt-4o-mini 32.75 27.86 28.01 13.67 12.32 12.10 316
gpt-4o-mini w/ Elab. 28.54 28.67 26.51 13.63 14.11 12.81 392
Llama-3.1-8b-Instruct 16.72 47.28 22.53 6.87 18.72 9.14 1282
Llama-3.1-8b-Instruct w/ Elab. 15.98 46.33 21.73 7.11 20.02 9.58 1253
Llama-3.1-70b-Instruct 21.70 48.09 27.97 8.93 20.26 11.59 857
Llama-3.1-70b-Instruct w/ Elab. 16.72 39.83 22.06 7.35 16.76 9.52 959
distilroberta-base ($k=3$, finetuned) 14.52 26.32 16.37 5.93 10.12 6.57 -
distilroberta-base ($k=5$, finetuned) 13.03 26.14 15.52 6.38 11.84 7.43 -
Table 6: Micro-averaged precision, recall, and F1 scores of different models on detecting friction. Each conversation contributes equally to the final metrics regardless of the number of friction instances it contains. #Predictions refer to the total number of instances of conversational friction found by each model. For reference, annotators identified 238 instances in total. gpt-4o without Elaboration of technical terms (Sec 3.1) performed best across all models. Macro-averaged results can be found in Table 11 in the Appendix.
@@ 5.1 Experimental Setup
Before moving on to prompting, we first explore a baseline setting in which we finetune a small encoder-based model distilroberta-base Sanh et al. (2019) on five randomly-split folds of Ubuntu-CG (Appendix A.2). Given an excerpt of a conversation consisting of a target turn $t$ and a context window of $k$ turns before and after $t$, the model is trained to predict whether $t$ is part of an annotated instance of friction. We report results on context windows of three and five, and notice no significant improvement in using a higher $k$. Finetuning a lighter-weight model enables us to understand the extent to which friction can be identified from surface features.
@@
Given a full conversation, we prompt several larger decoder-only LLMs to output a list of turn intervals exhibiting friction (Prompt A.4 in Appendix). [^12] We experimented with several prompting strategies such as adding random exemplars, self-consistency, and chain-of-thought reasoning, but found that they did not beat the F1 scores obtained simply by asking the model to detect friction windows along with brief explanations of why a dialog window represents friction. For each predicted turn interval, the model must provide a brief explanation for the cause of friction. We also include a setting where we provide models with the elaborations of technical terminology as outlined in §3.1 (“w Elab”). Note that in this setting, an LLM predicts possible friction intervals in a single pass, in contrast to the encoder baseline, where models make predictions on every single turn separately, taking neighboring turns as context.
@@
We evaluate models in the Friction Found and Friction Overlap settings 3.2. While Friction Found allows models like Llama-3.1-8b-Instruct Touvron et al. (2023) to obtain high recall scores by over-predicting friction intervals, Friction Overlap penalizes this behavior. For all experimental settings, we set temperature to 0.010.010.01.
@@ 5.2 Results
The F1 scores of our baseline encoder-only models (Table 6) in two settings (context window $k=3$ and $k=5$) give us an estimate of the degree to which friction is identifiable from shallower, local features as opposed to more complex, contextual, and implicit cases of friction like subtle clarification questions or other pragmatic phenomena. Error analysis reveals that instances predicted correctly by our baselines often contained explicit markers of friction, such as a user expressing frustration or dissatisfaction. While these baselines were competitive with some decoder-only models, especially under macro-averaging (Table 11 in Appendix) which weights individual friction instances equally regardless of conversation, the lack of model explanations limits deeper interpretation beyond error analysis. However, it does indicate that our dataset contains learnable surface features such as explicit expressions of frustration or anger that can be captured by a small encoder-only model.
Under both evaluation settings, gpt-4o without any further technical elaborations obtained the highest F1 score. This setting is used for all further error analysis and ablations. Models over-predict friction intervals (see #Predictions in Table 6).
@@
Explaining technical terms with gpt-4o helped our human annotators better understand the flow of information in a conversation. However, in the relaxed evaluation setting of (Friction Found), adding elaborations does not improve prediction scores of models. For Llama-3.1-8b-Instruct and gpt-4o-mini, adding elaborations improves recall and hence the overall F1 score. This may be due to elaborations “sharpening” the predicted intervals.
@@
We evaluate the binary prediction task with Cohen’s $\kappa$, framing it as inter-rater agreement between models and humans. Models’ over-prediction of friction intervals persists in the conversation level as well (Table 7). Predictions on task success, on the other hand, is highly correlated with annotator ratings of success.
@@ 6 Error Analysis
We now investigate the successes and failures of gpt-4o, the strongest performing model at this task.
@@
As a conversation proceeds, detecting friction requires a deeper understanding of preceding turns. To explore whether the position of friction impacts model accuracy, we stratify our results by conversational depth and calculate the relative depth of each instance of friction as the ratio of the first turn of the friction interval to the conversation length multiplied by 100. The mean relative depth of a detected instance of friction (35.1935.1935.19) is significantly smaller than the mean relative depth of a detected instance (49.6249.6249.62), according to an independent t-test ($p<0.01$). This indicates that models struggle with taking a longer context into account while determining whether participants’ versions of common ground are misaligned.
@@
Models, particularly gpt-4o, are more likely to correctly identify friction when an explicit request for conversational repair is present. Specifically, 77.22% of detected frictions involved an explicit RequestRepair, compared to 64.8164.8164.81% of frictions that went undetected ($p<0.05$). This highlights the tendency of models to rely on overt cues that signal a common ground misalignment.
Consider the conversation in Table 8. A’s utterance “how about nmap” (Turn 22) is not introducing nmap as an option, but following up on B’s earlier suggestion in Turn 21 by asking how nmap can be used to solve the issue. B reveals that they did not understand this interpretation through their response in Turn 23 (“yeah, i said nmap,”), prompting A to issue a Repair act. We hypothesize that this unconventional way of issuing a Repair (through a question) without an explicit RequestRepair results in an undetected conversational friction.
@@ 24 A (asker) I mean how do I use nmap to find that out?
Most model explanations accurately captured the cause of friction, with 57.8157.8157.81% instances annotated as equivalent and 34.3734.3734.37% as somewhat similar. Only 7.87.87.8% explanations did not point out the cause of friction at all. Echoing our earlier findings, models struggle to pinpoint the cause of friction even when they identified the window correctly. For example, in Figure 2, Row 1, A mistakenly assumes that B’s suggestion for a command (Turn 25) includes the keyword “try”, leading to the error “try command not found” (Turn 32) after which B repeats the dmesg command removing “try” at the beginning. LLMs must be able to pinpoint causes of friction to issue repairs that address the friction directly, a crucial ability in settings where LLMs are used for dispute resolution Tan et al. (2024).